Please indicate the correct answer and why.Thank you!Jose installs home security systems His fixed costs are $1000 a month and it costs him $24 of labor to install one system. The table shows the demand schedule for Joses installation services Price (dollars per installation) 40 32 24 16 Quantity (installations per month) 16 32 48 64 80 Calculate Joses profit-maximizing output, price, and economic profit. Joses profit-maximizing output isinstallations a month Joses profit-maximizing price is sDen installation. Joses is >>> If the firm incurs an economic loss, select loss in the dropdown box and do not enter a minus sign. of S a month.
Answer #1

4.4-4 4 4 4 22.2222 028082 32-30 02-8 8 5.7 75 0-16 32 48 64 40 32 16

Since the profit-maximizing condition are

MR=MC

or MR>MC

Hence according to the second condition the profit-maximizing price is =$32

Quantity=16 units

Profit=TR-TC

=512- (FC+VC)

=512- (1000+24*16)

=512- (1384)

=-($872)

Loss= $872

Hence Jose's loss is =$872 per month.

0